OpenStudy (anonymous):

Determine two pairs of polar coordinates for the point (3, -3) with 0° ≤ θ < 360°

OpenStudy (anonymous):

x = radius * cos(theta) y = radius * sin(theta) x = r*cos(theta) y = r*sin(theta) x^2 + y^2 = r^2, find r then plug in given x,y and r into the equations above to find theta, then you can express in polar coords, one will have a negative radius the other with have a positive radius

OpenStudy (anonymous):

by negative radius i mean negative r

OpenStudy (anonymous):

radial coordinate*